Write-Right, a vertically integrated firm produces both paper and writing tablets. The demand for tables is given by PT = 1.00 – 0.001Q where Q is the quantity of tablets. The marginal cost of producing the paper necessary for tablet is MCP = 0.20 + 0.001Q It costs the firm $0.10 to make the paper into a writing tablet. If there is no external market for the paper, what transfer price should top management set for the paper?
1
Expert's answer
2018-02-21T08:55:07-0500
1) So, if MCP = 0.20 + 0.001Q and PT = 1.00 – 0.001Q, MCP = PT,
0.20 + 0.001Q = 1.00 – 0.001Q
Q = 800, and so PT = 1.00 – 0.001Q = 0.2